# Attention as an Adaptive Filter

### v1

@@ B.3 Normal plus Rank-1 Noise Covariance
Recall that:
@@
Let us now assume that $\mathbf{Q}$ is normal plus rank-1, i.e. $\mathbf{Q}=\mathbf{S}\left(\bm{\Lambda}_{Q}+\mathbf{q}\mathbf{q}^{\dagger}\right)\mathbf{S}^{\dagger}$. We already derived the diagonal part. The rank-1 part is:
$\int_{0}^{|\Delta t_{ij}|}e^{-\bm{\Lambda}\tau}\big{(}\mathbf{q}\mathbf{q}^{\dagger}\big{)}e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$
$\int_{0}^{|\Delta t_{ij}|}\big{(}e^{-\bm{\Lambda}\tau}\mathbf{q}\big{)}\big{(}e^{-\bm{\Lambda}\tau}\mathbf{q}\big{)}^{\dagger}d\tau$
If we take: $\mathbf{q}(t_{j}+\tau)=e^{\bm{\Lambda}(\Delta t_{ij}+\tau)}\mathbf{p}$, so that $\mathbf{q}$ moves with the LTI’s flow, this becomes:
$e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}d\tau=e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}|\Delta t_{ij}|$
@@
$\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\big{(}\mathbf{q}\mathbf{q}^{\dagger}\big{)}e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}d\tau$
@@
$ae^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}d\tau=a\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}e^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
@@
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\sigma^{2}\kappa_{B}(\Delta t_{ij})\Big{(}\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big{)}$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big{[}\lambda_{C}^{2}\sigma^{2}\kappa_{B}(|\Delta t_{ij}|)\Big{(}\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big{)}+\eta^{2}\Big{(}\mathbf{I}+be^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}\mathbf{p}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big{)}\Big{]}e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\lambda_{C}^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)\Big{(}\mathbf{I}+a\mathbf{p}\mathbf{p}^{\dagger}\Big{)}+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big{(}\mathbf{I}+b\mathbf{p}\mathbf{p}^{\dagger}\Big{)}$
$\Big{(}\lambda_{C}^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{I}+\Big{(}a\lambda_{C}^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+b\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{p}\mathbf{p}^{\dagger}$
In particular, letting $\mathbf{p}=\mathbf{u}_{\mathbf{s}i}$, and setting $\sigma=\sigma_{t}$, $\eta=\eta_{t}$, $a=\frac{\sigma_{r}^{2}-\sigma_{t}^{2}}{\sigma_{t}^{2}}$, $b=\frac{\eta_{r}^{2}-\eta_{t}^{2}}{\eta_{t}^{2}}$, this becomes:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\Big{(}\lambda_{C}^{2}\sigma_{t}^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta_{t}^{2}e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{I}+\Big{(}\lambda_{C}^{2}(\sigma_{r}^{2}-\eta_{t}^{2})\kappa_{F}(|\Delta t_{ij}|)+(\eta_{r}^{2}-\eta_{t}^{2})e^{-2\alpha\Delta t_{ij}}\Big{)}\mathbf{u}_{\mathbf{s}i}\mathbf{u}_{\mathbf{s}i}^{\dagger}$
@@
Hence, the normal plus rank-1 process noise covariance model reduces exactly to the RT-SDE case: while $\mathbf{p}$ is not restricted to be a unit vector, any scaling of $\mathbf{p}$ can be absorbed into the radial noise parameters, so the model introduces no additional degrees of freedom beyond RT-SDE. The requirement that $\mathbf{p}$ flow with the LTI dynamics (which is necessary for the integral to be tractable), with $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, constrains the rank-1 component to rotate with the system, forcing the normal-plus-rank-1 noise model to align exactly with the RT-SDE structure.
B.4 Expressivity of Radial-Tangential SDEs
Proposition 1: Expressivity of Deterministic RT-SDEs.
@@ B.4 Expressivity of Radial-Tangential SDEs
$\left\|e^{\mathbf{A}t_{k}}-\prod_{\ell=1}^{L}e^{\mathbf{A}_{\ell}\Delta t}\right\|<\epsilon,\quad\forall k=1,\dots,m.$
Proof Sketch.
@@
Proposition 2: Expressivity of RT-SDE Covariance.
@@
$\big{\|}\bm{\Sigma}(t_{i})-\bm{\Sigma}^{\star}\big{\|}<\varepsilon,$

### v2

@@ B.6.1 Normal plus Rank-1 Noise Covariance
Recall that the propagated covariance in the eigenbasis is:
@@
Let us now assume that $\mathbf{Q}$ is normal plus rank-1, i.e. $\mathbf{Q}=\mathbf{S}\left(\bm{\Lambda}_{Q}+\mathbf{q}_{ij}(\tau)\mathbf{q}_{ij}^{\dagger}(\tau)\right)\mathbf{S}^{\dagger}$, where $\mathbf{q}_{ij}(\tau)$ may depend on $\Delta t_{ij}$.
We already derived the diagonal part. The rank-1 part is:
$\int_{0}^{|\Delta t_{ij}|}e^{-\bm{\Lambda}\tau}\big(\mathbf{q}_{ij}(\tau)\mathbf{q}_{ij}^{\dagger}(\tau)\big)e^{-\bm{\Lambda}^{\dagger}\tau}d\tau$
$\int_{0}^{|\Delta t_{ij}|}\big(e^{-\bm{\Lambda}\tau}\mathbf{q}_{ij}(\tau)\big)\big(e^{-\bm{\Lambda}\tau}\mathbf{q}_{ij}(\tau)\big)^{\dagger}d\tau$
If we take: $\mathbf{q}(t_{j}+\tau)=e^{\bm{\Lambda}(\Delta t_{ij}+\tau)}\mathbf{p}_{i}$, so that $\mathbf{q}_{ij}$ moves with the LTI’s flow, this becomes:
$e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}d\tau=e^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}|\Delta t_{ij}|$
@@
$\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}e^{-\bm{\Lambda}_{\Omega}\tau}\big(\mathbf{q}_{ij}\mathbf{q}_{ij}^{\dagger}\big)e^{-\bm{\Lambda}_{\Omega}^{\dagger}\tau}d\tau$
@@
$ae^{\bm{\Lambda}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}^{\dagger}\Delta t_{ij}}\int_{0}^{|\Delta t_{ij}|}e^{-2\alpha\tau}d\tau=a\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}e^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}$
@@
$\bm{\Lambda}_{\mathbf{V}B}(t_{i},t_{j})=\sigma^{2}\kappa_{B}(\Delta t_{ij})\Big(\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big)$
@@
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=e^{-\bm{\Lambda}\Delta t_{ij}}\Big[\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{B}(|\Delta t_{ij}|)\Big(\mathbf{I}+ae^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big)$
$+\eta^{2}\Big(\mathbf{I}+be^{\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}e^{\bm{\Lambda}_{\Omega}^{\dagger}\Delta t_{ij}}\Big)\Big]e^{-\bm{\Lambda}^{\dagger}\Delta t_{ij}}$
$\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)\Big(\mathbf{I}+a\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}\Big)+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big(\mathbf{I}+b\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}\Big)$
$\Big(\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{I}+\Big(a\big|\lambda_{C}\big|^{2}\sigma^{2}\kappa_{F}(|\Delta t_{ij}|)+b\eta^{2}e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{p}_{i}\mathbf{p}_{i}^{\dagger}$
Letting $\mathbf{p}_{i}=\mathbf{u}_{\mathbf{s}i}$, and setting $\sigma=\sigma_{t}$, $\eta=\eta_{t}$, $a=\frac{\sigma_{r}^{2}-\sigma_{t}^{2}}{\sigma_{t}^{2}}$, $b=\frac{\eta_{r}^{2}-\eta_{t}^{2}}{\eta_{t}^{2}}$, this reduces to the RT-SDE case:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})$ $=\Big(\big|\lambda_{C}\big|^{2}\sigma_{t}^{2}\kappa_{F}(|\Delta t_{ij}|)+\eta_{t}^{2}e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{I}$
$+\Big(\big|\lambda_{C}\big|^{2}(\sigma_{r}^{2}-\eta_{t}^{2})\kappa_{F}(|\Delta t_{ij}|)+(\eta_{r}^{2}-\eta_{t}^{2})e^{-2\alpha\Delta t_{ij}}\Big)\mathbf{u}_{\mathbf{s}i}\mathbf{u}_{\mathbf{s}i}^{\dagger}$
@@
In summary, the co-rotating normal-plus-rank-1 noise model provides a minor generalization of the RT-SDE structure by permitting anisotropy along directions not tied to the state. However, this comes at the cost of losing the clean hyper-spherical MLE formulation.
B.6.2 Expressiveness of Radial-Tangential SDEs
The efficiency of the memory-optimized AFA comes from restricting the dynamics to isotropic decay and noise. We now examine the expressive power of these constraints. Proposition 1 shows that a sequence of linear systems with isotropic decay can approximate the flow of any diagonalizable LTI system to arbitrary accuracy. Proposition 2 shows that a multi-head Radial-Tangential SDE can approximate any target covariance at a fixed time, and thus represent arbitrary state-dependent uncertainty structures.
Proposition 1: Expressiveness of Deterministic RT-SDEs.
@@ B.6.2 Expressiveness of Radial-Tangential SDEs
$\Big|e^{\mathbf{A}t_{k}}-\prod_{\ell=1}^{L}e^{\mathbf{A}_{\ell}\Delta t}\Big|<\epsilon,\quad\forall k=1,\dots,m.$
Proof Sketch:
@@
Proposition 2: Expressiveness of RT-SDE Covariance.
@@
$\big|\bm{\Sigma}(t_{i})-\bm{\Sigma}^{\star}\big|<\varepsilon,$